# Usman Ahmed,a Mikael P. Johansson,a,b Susi Lehtola,a and Dage Sundholm∗,a

### v1

@@ 3.6 GGA functionals
The revPBE-NL and XLYP-D3 functionals are the best performing functionals among the GGAs. Their performance drops from the 14th place to 124th, and 17th rank to 141th without dispersion corrections, respectively. The rPBE-D4 and BLYP-D3 functionals are the next ones on the GGA ranking list ranked 19th and 21st, respectively. The BOP functional places 23rd. Other GGA functionals among the first 50 on the ranking list are PBE-D3 and PW91-D3, which are ranked 48th and 49th, respectively. The rest of the GGA functionals including dispersion effects appear among the second third on the ranking lists. The ranking of HF is 145th, which improves to 46th and 107th with D3 and D4 dispersion corrections, respectively.
@@ 4 Summary and conclusions
We calculated the hydrogen-bonding strength for 14 molecules using 152 density functional approximations (DFAs). The accuracy of the calculated bond strengths with various DFAs was assessed by comparison to the literature CCSD(T)-cf energies of Ahmed et al.1, which they extrapolated to the complete basis set limit, as well as to the Schrödinger limit using the continued-fraction methodology of Goodson51.
We began the work by validating the numerical reliability of our computational method by assessing the convergence of the hydrogen-bonding strengths across the Karlsruhe basis set family for the B97M-D3 functional. Taking the def2-QZVPPD values as reference, we observed that diffuse functions are important to obtain converged interaction energies, and we therefore recommend their use in modeling hydrogen bonds. As the def2-TZVPD and def2-TZVPPD values are within 0.5 kJ/mol of the def2-QZVPPD data, we can safely assume that the quadruple-$\zeta$ reference calculations are indeed of benchmark quality. The def2-QZVPPD basis set is often used for benchmarking in the literature, and has been employed in the fitting of the $\omega$B97M-V functional,138 for example.
Having validated the numerical methodology, we continued with the benchmark. We found the top of the list to be dominated by the Berkeley functionals, either with or without dispersion. The most accurate functional for hydrogen-bonding strengths was found to be B97M-D3, which is the B97M-V functional whose VV10 dispersion correction has been replaced by the empirical D3 dispersion correction. We found the accuracy of mGGA functionals to be surprisingly sensitive to the description of the dispersion energy, tentatively supporting the recent findings of Montero de Hijes et al.24. Other functionals that accurately predicted hydrogen-bonding strengths included the range-separated $\omega$B97M-V mGGA functional, the $\omega$B97, $\omega$B97X, $\omega$B97X-V range-separated GGA functionals, as well as the semi-local M11-L-D3 mGGA and the range-separated M11-D3 mGGA functional.

### v2

@@ 3.5 Global hybrid functionals
MAX values generally increase with increasing MAE value. However, revPBE0-D3 and M06-2X have much smaller MAX values than some of the higher ranked DFAs. Judged from the MAX value, revPBE0-D3 and M06-2X would be ranked 12th and 17th instead of 28th and 29th based on the MAE value, respectively.
@@ 3.6 GGA functionals
The revPBE-NL and XLYP-D3 functionals are the best performing DFAs among the GGAs. Their performance drops from the 14th place to 124th, and 17th rank to 141th without dispersion corrections, respectively. The rPBE-D4 and BLYP-D3 functionals are the next ones on the GGA ranking list ranked 19th and 21st, respectively. The BOP functional places 23rd. Other GGA functionals among the first 50 on the ranking list are PBE-D3 and PW91-D3, which are ranked 48th and 49th, respectively. The rest of the GGA functionals including dispersion effects appear among the second third on the ranking lists. The ranking of HF is 145th, which improves to 46th and 107th with D3 and D4 dispersion corrections, respectively.
@@ 4 Summary and conclusions
We calculated the hydrogen-bonding strength for 14 dimers using 152 density functional approximations (DFAs). The accuracy of the calculated bond strengths with various DFAs was assessed by comparison to the literature CCSD(T)-cf energies of Ahmed et al.1, which they extrapolated to the complete basis set limit, as well as to the Schrödinger limit using the continued-fraction methodology of Goodson51.
We began the work by validating the numerical reliability of our computational method by assessing the convergence of the hydrogen-bonding strengths across the Karlsruhe basis set family for the B97M-D3 functional. Taking the def2-QZVPPD values as reference, we observed that diffuse functions are important to obtain converged interaction energies, and we therefore recommend their use in modeling hydrogen bonds. As the def2-TZVPD and def2-TZVPPD values are within 0.5 kJ/mol of the def2-QZVPPD data, we can safely assume that the quadruple-$\zeta$ reference calculations are indeed of benchmark quality. The def2-QZVPPD basis set is often used for benchmarking in the literature, and has been employed in the fitting of the $\omega$B97M-V functional,141 for example.
Having validated the numerical methodology, we continued with the benchmark. We found the top of the list to be dominated by the Berkeley functionals, either with or without dispersion. The most accurate DFA for hydrogen-bonding strengths was found to be B97M-D3, which is the B97M-V functional whose VV10 dispersion correction has been replaced by the empirical D3 dispersion correction. We found the accuracy of mGGA functionals to be surprisingly sensitive to the description of the dispersion energy, tentatively supporting the recent findings of Montero de Hijes et al.24. Other DFAs that accurately predicted hydrogen-bonding strengths included the range-separated $\omega$B97M-V mGGA functional, the $\omega$B97, $\omega$B97X, $\omega$B97X-V range-separated GGA functionals, as well as the semi-local M11-L-D3 mGGA and the range-separated M11-D3 mGGA functional.
@@ Acknowledgement
  • 52 Jmol: an open-source Java viewer for chemical structures in 3D. http://www.jmol.org.
@@
  • Jansen and Ros 1969 H. Jansen and P. Ros, Non-empirical molecular orbital calculations on the protonation of carbon monoxide, Chem. Phys. Lett., 1969, 3, 140–143, DOI: 10.1016/0009-2614(69)80118-1.
  • Boys and Bernardi 1970 S. F. Boys and F. Bernardi, The calculation of small molecular interactions by the differences of separate total energies. Some procedures with reduced errors, Mol. Phys., 1970, 19, 553–566, DOI: 10.1080/00268977000101561.